OpenStudy (anonymous):

04.03 Exponential Functions and Models!!! someone help :(?? MEDAL?!

OpenStudy (anonymous):

Today is your lucky day! You have just won a contest and they can pay you your prize in two different ways. Option one is to pay you $10,000 every ten days, for thirty days. Option two is a prize of a penny that doubles every day, for thirty days.

OpenStudy (anonymous):

@noneyabusiness ???

OpenStudy (anonymous):

@.WhiteDragon. ??

OpenStudy (anonymous):

@thewonderfuladele??

OpenStudy (anonymous):

If you selected this option, on day 30 you will receive $10,737,418.24! This is an example of an exponential function because the rate of increase is changing. It gets twice as big each day! OPTION 2

OpenStudy (anonymous):

If you selected this option, you will have $30,000 at the end of the thirty days. This is an example of a linear function because the rate of increase is constant. OPTION 1

OpenStudy (anonymous):

@thewonderfuladele please help?

OpenStudy (anonymous):

@srossd??

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@CFKane

OpenStudy (anonymous):

@texaschic101?

OpenStudy (anonymous):

@jim_thompson5910??

jimthompson5910 (jim_thompson5910):

What's your question? I see the problem laid out, but I'm not sure where you're stuck.

OpenStudy (anonymous):

I am just stuck period... I can take a picture of the problem?

jimthompson5910 (jim_thompson5910):

alright, go ahead

OpenStudy (anonymous):

okay one second.

OpenStudy (anonymous):

04.03 Assignment Iris has been studying an invasive population of snails. This particular snail has no local predators so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

OpenStudy (anonymous):

I got the real questions..

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

were you able to get anywhere with these new questions?

OpenStudy (anonymous):

no :(

jimthompson5910 (jim_thompson5910):

it says to "Create your own exponential function", so are you given a data set to work with?

OpenStudy (anonymous):

No sir..

jimthompson5910 (jim_thompson5910):

odd how they expect you to create a function, but don't give you any data

jimthompson5910 (jim_thompson5910):

it's like they want you to make up a function out of the blue, hmm

OpenStudy (anonymous):

exactly.

jimthompson5910 (jim_thompson5910):

does it give you a growth rate and initial population?

OpenStudy (anonymous):

No :(

jimthompson5910 (jim_thompson5910):

ok there's a lot of missing information for the first question

jimthompson5910 (jim_thompson5910):

it cannot be answered unless you want to make something up but that may be wrong

OpenStudy (anonymous):

A local snail population grows according to the function g(x) = 200(1.03)2x. in question 2

jimthompson5910 (jim_thompson5910):

A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

jimthompson5910 (jim_thompson5910):

hint: \[\Large a^{b*c} = \left(a^b\right)^c\]

OpenStudy (anonymous):

I still do not get it :(....

jimthompson5910 (jim_thompson5910):

for example \[\Large 2^{3x} = (2^3)^x\]

jimthompson5910 (jim_thompson5910):

what is 2^3?

OpenStudy (anonymous):

2*2*2=6

jimthompson5910 (jim_thompson5910):

more like 2*2*2 = 8

jimthompson5910 (jim_thompson5910):

* means multiply

jimthompson5910 (jim_thompson5910):

so, \[\Large 2^{3x} = (2^3)^x = 8^x\]

jimthompson5910 (jim_thompson5910):

Notice how I went from the exponent of 2x to just 'x'

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so how can we apply that to \[\Large 1.03^{2x}\]

OpenStudy (anonymous):

mulitpy 8 and 2?

jimthompson5910 (jim_thompson5910):

no, \[\Large 1.03^{2x} = (1.03^{2})^{x} = ???\]

OpenStudy (anonymous):

I am still completely lost.. :(

OpenStudy (anonymous):

I need help on this too!

OpenStudy (anonymous):

What would be a good domain/range?

OpenStudy (anonymous):

was this assignment ever done ?